# MoLA: Motion Generation and Editing with Latent Diffusion Enhanced by Adversarial Training

### v2

@@ 4.4. Motion Generation
Next, we evaluate inference cost and quality of generation in comparison to existing text-to-motion methods that have publicly available implementations (Tevet et al., 2023; Chen et al., 2023; Zhang et al., 2023b; Pinyoanuntapong et al., 2024; Guo et al., 2023). Since our requirements are 1) fast generation, 2) high generation performance, and 3) training-free editing, we apply adversarial training to a continuous latent case, specifically MLD (Chen et al., 2023) (introduced in Section 3) We show that our proposed strategy improves MLD and achieves a better performance in Figure 3. Note that the difference between the MLD baseline and our MoLA is the training method, not the inference. Thus, as shown in Figure 3, inference with our model can be achieved with equivalent time to the MLD baseline, which is much faster in generation than data-space diffusion MDM (Tevet et al., 2023) that is the only existing method that can perform training-free motion editing. It is true that most methods based on discrete representation perform very well in terms of fidelity and speed, but there is no discrete method that can deal with multiple editing tasks in a training-free manner and satisfies our requirements.
@@
In addition, we evaluate our proposed MoLA in comparison to current SOTA methods (Zhang et al., 2024; Tevet et al., 2023; Wang et al., 2023; Zhong et al., 2023; Guo et al., 2023; Lou et al., 2023; Pinyoanuntapong et al., 2024; Zhang et al., 2023b; Chen et al., 2023) using five metrics (R-Precision, Fréchet Inception Distance (FID), Mulit-modal distance (MMDist), Diversity, and MultiModality (MModality)) proposed by Guo et al. (Guo et al., 2022). For evaluation, we select the model that achieves the best FID, which is a metric that evaluates the overall motion quality, on the validation set and report its performance on the test set on each of the two datasets. We show the results in Tables 4 and 5. The methods are organized into three groups: i) those using VQ-based latent representations (Discrete), ii) those using data-space diffusion model (Continuous (raw data)), and iii) those using VAE-based latent representations (Continuous (latent)). The discrete approaches perform well in motion generation (e.g., (Guo et al., 2023; Lou et al., 2023; Pinyoanuntapong et al., 2024)). However, those models cannot control an arbitrary set of joints in a training-free manner (Pinyoanuntapong et al., 2024) as we have discussed so far. MDM (Tevet et al., 2023), MotionDiffuse (Zhang et al., 2024) and Fg-T2M (Wang et al., 2023) adopt data-space diffusion, while MLD (Chen et al., 2023) and our MoLA utilize a diffusion model in a low-dimensional latent space. Therefore, the former are grouped in Continuous (raw data) and the latter in Continuous (latent) in the tables. As shown in Figure 3, MLD and our MoLA can be much faster in generation than MDM, which is the only existing method that can perform training-free motion editing. Moreover, MoLA achieves a better performance than MLD, especially in the R-precision, FID, and MMDist metrics on both HumanML3D and KIT-ML datasets as shown in the tables. As we will show in the next section, our MoLA can also deal with multiple motion editing tasks without additional training, which means our MoLA satisfies our three requirements best.
Top-1 Top-2 Top-3
MoMask (Guo et al., 2023) $0.433^{\pm.007}$ $0.656^{\pm.005}$ $0.781^{\pm.005}$ $0.204^{\pm.011}$ $2.779^{\pm.022}$ $1.131^{\pm.043}$
MMM (Pinyoanuntapong et al., 2024) $0.381^{\pm.005}$ $0.590^{\pm.006}$ $0.718^{\pm.005}$ $0.429^{\pm.019}$ $3.146^{\pm.019}$ $10.633^{\pm.097}$ $1.105^{\pm.026}$
Continuous (latent) MLD (Chen et al., 2023) $0.390^{\pm.008}$ $0.609^{\pm.008}$ $0.734^{\pm.007}$ $0.404^{\pm.027}$ $3.204^{\pm.027}$ $10.80^{\pm.117}$ $2.192^{\pm.071}$
MoLA (ours) $0.401^{\pm.004}$ $0.612^{\pm.004}$ $0.728^{\pm.004}$ $0.350^{\pm.020}$ $3.187^{\pm.027}$ $10.88^{\pm.110}$ $2.226^{\pm.097}$
Table 5. Comparison with state-of-the-art methods on KIT-ML.
@@ 4.5. Motion Editing
Here, we demonstrate three different types of editing tasks with a single framework: editing in the time direction (in-betweening), editing in some joints of motion (upper body editing), and editing with the path specified as the trajectory (path following). We experiment with the following motion editing on models trained in HumanML3D (Guo et al., 2022).
Figure 4. Qualitative results of motion in-betweening task. Blue and red points are used as inputs to constrain the start and end positions, respectively. These are the results generated by our framework from the same text prompt (“A person walks while raising his left hand.”) for two different start-end positions.
In-betweening: Motion in-betweening is an important editing task that interpolates or fills the gaps between keyframes or major motion joints to create smooth 3D motion animation. In our framework, by applying the update rule of MPGD (He et al., 2024) on latent space (Section 3.4), it is easy to coordinate and generate motion between past and future contexts without additional training. We only need to set the start-end positions or the motions of a few frames as the control signal $\bm{y}$ in (12). The motion in-betweening results with our model when different start-end controls are given as $\bm{y}$ in the same text condition are shown in Figure 4.
Figure 5. Qualitative results of upper body editing task. Each colored point is used inputs to control the lower body position as a constraint. These are the results generated by our framework from the same text prompt (“A person puts his right hand to the air.”) for two different lower body positions.
Upper body editing: Upper body editing combines generated upper body parts with given lower body parts. Generating certain joints to fit given body parts can be seen as the task of outpainting in the spatial dimension of motion. In our framework, similar to the motion in-betweening, the loss and its gradient are calculated in the data-space as (11), and sampling is performed taking into account its gradients. The control signal $\bm{y}$ in (12) is set as lower body positions that is not subject to editing. Note that although we are dealing with upper body editing in these experiments, it is in principle possible to specify a different joint subset. Therefore, without the need for a separate architecture or additional training, our framework allows editing conditioned on given body parts of motion. The upper body editing results with our model when different lower body controls are given as $\bm{y}$ in the same text condition are shown in Figure 5.
Figure 6. Qualitative results of path-following task. The points are the input to control a motion trajectory. These are the results generated by our framework from the same text prompt (“A person walks with his hands up.”) for two different given trajectories.
@@
Path-following: Path-following motion generation is a task of giving a trajectory (often the position of the pelvis) and generating the motion that matches with the given route. Controlling the trajectory of generated motion can lead to more motion variation, avoidance of obstacles, and the creation of motion that meets physical constraints. Our framework can edit a motion to follow a specific path by minimizing the loss between a joint of the generated motion and a desired trajectory of the joint, similar to motion in-betweening. In this experimental case, the desired pelvis trajectory is set as the control signal $\bm{y}$ in (12). The path-following generation results with our model when different path controls are given as $\bm{y}$ in the same text condition are shown in Figure 6.
@@ 5. Conclusion
In this paper, we proposed MoLA, a text-to-motion model that achieves fast, high-quality generation with multiple control tasks in a single framework. To construct our model, we employed a latent diffusion-based method, and introduced adversarial training. In addition, we applied a training-free guided generation framework based on MPGD. The results of quantitative evaluation demonstrate the effectiveness of adversarial training in the text-to-motion generation. In particular, we found that replacing GAN with SAN boosts the performance of our model. We also showed that our model can perform motion in-betweening, upper body editing, and path following within a single framework.
Appendix A Notation
In Table 6, we list the several notations and symbols that we have used in this paper.
Notation/Symbol Description
$\mathbb{N}$ The sets of natural numbers
$\mathbb{R}$ The sets of real numbers
$\mathbb{R}_{+}$ The sets of positive real numbers
$N\in\mathbb{N}$ A motion feature dimension
$L\in\mathbb{N}$ A motion length
$d_{z}\in\mathbb{N}$ A latent space dimension
$d_{c}\in\mathbb{N}$ A conditional embedding dimension
$t\in\mathbb{N}$ A time step for diffusion model
$\bm{\eta}$ A trainable parameter for the encoder
$\bm{\psi}$ A trainable parameter for the decoder
$\bm{\phi}=\{\bm{\varphi},\bm{\omega}\}$ A trainable parameter for the discriminator
$\bm{\theta}$ A trainable parameter for the estimator in latent diffusion
$\bm{x}\in\mathbb{R}^{N\times L}$ A motion input
$\bm{z}\in\mathbb{R}^{d_{z}}$ A latent variable
$\tau(\bm{c})\in\mathbb{R}^{d_{c}}$ A output of text encoder corresponding to a prompt $\bm{c}$
$q_{\bm{\eta}}:\mathbb{R}^{N\times L}\to\mathbb{R}^{d_{z}}\times\mathbb{R}_{+}^{d_{z}}$ A encoder parameterized by $\bm{\eta}$ in Stage 1
$g_{\bm{\psi}}:\mathbb{R}^{d_{z}}\to\mathbb{R}^{N\times L}$ A decoder parameterized by $\bm{\psi}$ in Stage 1
$f_{\bm{\phi}}:\mathbb{R}^{N\times L}\to\mathbb{R}$ A discriminator parameterized by $\bm{\phi}$ in Stage 1
$\epsilon_{\bm{\theta}}:\mathbb{R}^{d_{z}}\times\mathbb{N}\times\mathbb{R}^{d_{c}}\to\mathbb{R}^{d_{z}}$ A conditional estimator parameterized by $\bm{\theta}$ in Stage 2
$\lambda_{\text{reg}}\in\mathbb{R}_{+}$ A weight of VAE regularization term in Stage 1
$\lambda_{\text{adv}}\in\mathbb{R}_{+}$ A weight of adversarial loss in Stage 1
$\alpha_{t}\in(0,1)$ A noise scheduling parameter in Stage 2
$\rho_{t}\in\mathbb{R}_{+}$ A time-dependent step size in guided generation
Table 6. Notation table
@@ Appendix B Editing procedure on guided generation
In Algorithm 1, we show the specific procedure for guided generation described in Section 3.4. Note that we apply a time-travel technique to the update rule in Section 3.4 as in (Lugmayr et al., 2022; Wang et al., 2022; Yu et al., 2023; He et al., 2024) to achieve better editing results. This technique adds noise after each gradient descent step to implicitly perform a multi-step optimization of minimizing the distance measuring function $L(\cdot,\bm{y})$ and lead to an improvement in the editing quality.
@@
0:  motion control signal $\bm{y}$, output of text encoder $\tau(\bm{c})$, estimator in latent diffusion $\epsilon_{\theta}(\cdot,t,\tau(\bm{c}))$, distance measuring function $L(\cdot;\bm{y})$, pre-defined parameter $\bar{\alpha}_{t}$, time-dependent step-size $\rho_{t}$, and the repeat times of time-travel of each step $\{r_{1},\cdots,r_{T}\}$.
@@
Methods FID $\downarrow$ R-Precision Top-1 $\uparrow$ MMDist $\downarrow$
T2M-GPT (Zhang et al., 2023b) $0.141^{\pm.005}$ $0.492^{\pm.003}$ $3.121^{\pm.009}$
T2M-GPT + GAN (ours) $0.136^{\pm.010}$ $0.485^{\pm.004}$ $3.116^{\pm.009}$
T2M-GPT + SAN (ours) $0.133^{\pm.006}$ $0.494^{\pm.003}$ $3.111^{\pm.013}$
MLD (Chen et al., 2023) $0.473^{\pm.013}$ $0.481^{\pm.003}$ $3.196^{\pm.010}$
MLD + GAN (ours) $0.371^{\pm.014}$ $0.482^{\pm.002}$ $3.173^{\pm.011}$
MLD + SAN (MoLA; ours) $0.362^{\pm.009}$ $0.514^{\pm.003}$ $3.162^{\pm.008}$
Table 7. Effectiveness of adversarial training for motion generation performance on HumanML3D
Methods FID $\downarrow$ R-Precision Top-1 $\uparrow$ MMDist $\downarrow$
T2M-GPT (Zhang et al., 2023b) $0.514^{\pm.029}$ $0.416^{\pm.006}$ $3.007^{\pm.029}$
T2M-GPT + GAN (ours) $0.484^{\pm.036}$ $0.408^{\pm.008}$ $3.027^{\pm.022}$
T2M-GPT + SAN (ours) $0.439^{\pm.033}$ $0.423^{\pm.006}$ $2.947^{\pm.030}$
MLD (Chen et al., 2023) $0.404^{\pm.027}$ $0.390^{\pm.008}$ $3.204^{\pm.027}$
MLD + GAN (ours) $0.382^{\pm.028}$ $0.385^{\pm.005}$ $3.305^{\pm.035}$
MLD + SAN (MoLA; ours) $0.350^{\pm.020}$ $0.401^{\pm.004}$ $3.187^{\pm.027}$
Table 8. Effectiveness of adversarial training for motion generation performance on KIT-ML
Appendix C Evaluation metrics details

### v3

@@ 4.1. Motion Generation
4.1.1. Performance comparison with other text-to-motion models
@@
We evaluate generation quality and inference cost in comparison to existing text-to-motion methods that have publicly available implementations (Tevet et al., 2023; Chen et al., 2023; Zhang et al., 2023b; Pinyoanuntapong et al., 2024b; Guo et al., 2024). As shown in Figure 2, inference with our model can be achieved with equivalent time to MLD, which is much faster in generation than data-space diffusion MDM (Tevet et al., 2023) that is the only existing method that can perform training-free motion editing. It is true that most methods based on discrete representation perform very well in terms of fidelity and speed, but there is no discrete method that can deal with multiple editing tasks in a training-free manner and satisfies our requirements.
@@
In addition, we evaluate our proposed MoLA in comparison to current SOTA methods (Zhang et al., 2024; Tevet et al., 2023; Wang et al., 2023; Zhong et al., 2023; Guo et al., 2024; Lou et al., 2023; Pinyoanuntapong et al., 2024b; Zhang et al., 2023b; Chen et al., 2023; Zou et al., 2024; Pinyoanuntapong et al., 2024a) using five metrics (R-Precision, Fréchet Inception Distance (FID), Multi-modal distance (MMDist), Diversity, and MultiModality (MModality)) proposed by Guo et al. (Guo et al., 2022). For evaluation, we select the model that achieves the best FID, which is a metric that evaluates the overall motion quality, on the validation set and report its performance on the test set of HumanML3D. We show the results in Table 1. The methods are organized into three groups: i) those using VQ-based latent representations (Discrete), ii) those using data-space diffusion model (Continuous (raw data)), and iii) those using VAE-based latent representations (Continuous (latent)). The discrete approaches perform well in motion generation (e.g., (Guo et al., 2024; Lou et al., 2023; Pinyoanuntapong et al., 2024b)). However, those models cannot control an arbitrary set of joints in a training-free manner (Pinyoanuntapong et al., 2024b) as we have discussed so far. MDM (Tevet et al., 2023), MotionDiffuse (Zhang et al., 2024) and Fg-T2M (Wang et al., 2023) adopt data-space diffusion, while MLD (Chen et al., 2023), MotionLCM (Dai et al., 2024) and our MoLA utilize a diffusion model in a lower-dimensional latent space. Therefore, the former are grouped in Continuous (raw data) and the latter in Continuous (latent) in the table. MoLA achieves the best performance in continuous methods, especially in the R-precision, FID, and MMDist metrics as shown in Table 1. Moreover, as shown in Figure 2, our MoLA are much faster in generation than MDM, which is the only existing method that can perform training-free motion editing.
Figure 4. Comparison of motion length distributions between the HumanML3D test set and the generated motion samples.
Method Reconstruction Generation
rFID $\downarrow$ MPJPE $\downarrow$ FID $\downarrow$ MMDist $\downarrow$
Dimension of latent space
MoLA ($d_{z}=8$) 0.110 54.2 0.183 3.099
MoLA ($d_{z}=16$) 0.030 29.3 0.115 3.008
MoLA ($d_{z}=32$) 0.028 26.8 0.904 3.536
Adversarial training
w/o GAN or SAN 0.038 29.3 0.126 3.053
w/ GAN instead of SAN 0.032 29.5 0.141 3.044
Input for the encoder $q_{\bm{\eta}}$
$[\bm{m}^{i},a]$ instead of Eq. (2) 0.029 31.2 0.112 3.024
Table 2. Analysis of motion reconstruction /generation performance on HumanML3D
Editing type Methods
R-Precision
Top-3 $\uparrow$
FID $\downarrow$ Diversity $\rightarrow$ Traj. err. $\downarrow$ Loc. err.$\downarrow$ Avg. err.$\downarrow$ AITS $\downarrow$
Overfitting-based editing OmniControl 0.688 0.192 9.533 0.065 0.007 0.053 74.4
MotionLCM 0.759 0.501 9.293 0.237 0.054 0.164 0.02
Training-free editing MoLA (ours) 0.761 0.486 9.322 0.271 0.051 0.159 1.04
Table 3. Comparison of motion control (path-following task) on HumanML3D dataset
Next, we discuss the effects of the activation variable in our motion representation, introduced in Section 3.1. By incorporating the activation variable into the motion features in Equation (2), our stage 2 model in Section 3.3 enables the generation of variable length motions. Figure 4 shows the distribution of sample lengths generated by our model compared to the test set of HumanML3D. It is demonstrated that our method can generate variable-length motions, whose distribution is close to that of real motions.
4.1.2. Ablation studies
Moreover, we discuss the effect of the dimension of latent space, adversarial training, and motion representation on the stage 1 model, and we utilize the reconstruction FID (rFID), MPJPE, FID, and MMDist as evaluation metrics. The dimensionality of the latent space may affect not only reconstruction quality but also influence the difficulty of training in stage 2, ultimately impacting the quality of the generated outputs. To investigate this, we evaluate the performance for cases with $d_{z}=\{8,16,32\}$. Table 2 shows that although the case of $d_{z}=16$ does not achieve the best generation quality compared to the other cases, it performs best in terms of FID and MMDist. Therefore, we adopted $d_{z}=16$ for MoLA. The results for VAE combined with GAN/SAN and modifying the encoder inputs are also shown in Table 2. As shown, the adversarial training is effective in the motion reconstruction task. It not only improves the reconstruction performance in stage 1 but also contributes to enhanced generation performance in stage 2. In particular, we can improve the performance of the stage 1 model by adopting the SAN framework instead of the conventional GAN. A better rFID is directly related to the upper bound of the overall performance of a text-to-motion model. Besides, the modifications to the encoder inputs in Equation (2) are shown to also have a comparable positive effect. Thus, we used VAE trained with the SAN framework (Section 3.2, Appendix) and the motion representation Equation (2) for our MoLA.
@@ 4.2. Motion Editing
Here, we demonstrate three different types of editing tasks with a single framework: editing with the path specified as the trajectory (path following), editing in the time direction (in-betweening), and editing in some joints of motion (upper body editing). In particular, for the path-following task, we quantitatively compare our approach with existing models (Xie et al., 2024; Dai et al., 2024).
@@
Path-following: Path-following motion generation is a task of giving a trajectory (often the position of the pelvis) and generating the motion that matches the given route. Controlling the trajectory of generated motion enables more motion variation, avoidance of obstacles, and the creation of motion that meets physical constraints. Our framework can edit a motion to follow a specific path by minimizing the loss between a joint of the generated motion and a desired trajectory of the joint, applying the update rule of MPGD (He et al., 2024) on latent space (Section 3.4). In this experimental case, the desired pelvis trajectory is set as the control signal $\bm{y}$ in Equation (13). Figure 5 shows the editing results with our model when different path controls are given as $\bm{y}$ in the same text condition.
In addition, we show a quantitative comparison with existing methods: OmniControl (Xie et al., 2024) and MotionLCM (Dai et al., 2024), where the former is Continuous (raw data) type method and the latter is Continuous (raw latent) one. We compare them in terms of FID, R-Precision, Diversity, Trajectory err (50cm), Location err (50cm), Avgerage err, and Average inference time per sentence (AIST) in Table 3, following (Dai et al., 2024). Table 3 demonstrates that MoLA achieves competitive editing performance across all compared to other editing methods. Notably, while MoLA shows lower performance in following control signals compared to OminiControl, it achieves significantly faster editing. Note that OmniControl and MotionLCM have been training for this specific editing task, while our MoLA performs path-following without model fine-tuning. In other words, MoLA can perform other tasks without additional training as shown in the following sections, but OmniControl and MotionLCM require training a separate model for each task. MoLA is a more flexible and efficient framework.
In-betweening: Motion in-betweening is an important editing task that interpolates or fills the gaps between keyframes or major motion joints to create smooth 3D motion animation. Our framework can coordinate and generate motion between past and future contexts without additional training. We only need to set the start-end positions or the motions of a few frames as the control signal $\bm{y}$ in Equation (13). Figure 6 depicts the motion in-betweening results with our model when different start-end controls are given as $\bm{y}$ in the same text condition.
Upper body editing: Upper body editing combines generated upper body parts with given lower body parts. Generating some joints while keeping the other body joints following a given control signal can be seen as the task of outpainting in the spatial dimension of motion. In our framework, the loss and its gradient are calculated in the data space as Equation (12), and sampling is performed taking into account its gradients. The control signal $\bm{y}$ in Equation (13) is set to lower body positions that are not subject to editing. Figure 7 shows the upper body editing results with our model when different lower body controls are given as $\bm{y}$ in the same text condition. Note that, although we are dealing with upper body editing in this experiment, it is in principle possible to specify a different joint subset. Therefore, without the need for a separate architecture or additional training, our framework allows editing conditioned on given body parts of motion.
@@ 5. Conclusion
In this paper, we proposed MoLA, a text-to-motion model that achieves fast, high-quality generation with multiple control tasks in a single framework. To construct our model, we rethought the representation of motion features and introduced an activation variable that characterizes the length of the motion. We also employed a latent diffusion-based method and introduced adversarial training. In addition, we applied a training-free guided generation framework based on MPGD. Our experiments showed that MoLA generates variable-length motions, whose distribution is close to that of real motions. Likewise, our quantitative evaluation demonstrated the effectiveness of adversarial training in text-to-motion generation. We also showed that our model can perform motion path-following, in-betweening, and upper body editing in a training-free manner.
Appendix A SAN-based discriminator
Takida et al. (Takida et al., 2024) incorporated a sliced optimal transport perspective into a general GAN and established a new framework, slicing adversarial network (SAN). Following this work, we first decompose the discriminator $f_{\bm{\phi}}$ into the last linear layer and the remaining neural part, denoted as $\bm{w}\in\mathbb{R}^{d_{w}}$ and $h_{\bm{\varphi}}:\mathbb{R}^{N\times L}\to\mathbb{R}^{d_{w}}$ with $d_{w}\in\mathbb{N}$, respectively. This decomposition provides an interpretation of the discriminator: the neural function $h_{\bm{\varphi}}$ maps motion sequences to non-linear features, and then the linear layer $\bm{w}$ projects them into scalars. As a preliminary step for applying SAN to the adversarial training in Section 3.2.1, we normalize $w$ using its norm, resulting in $\bm{\omega}=\bm{w}/\|\bm{w}\|_{2}$, which indicates the direction of the projecting. Now, the discriminator is represented in the form of an inner product as $f_{\bm{\phi}}(\bm{x})=\bm{\omega}^{\top}h_{\varphi}(\bm{x})$, and its parameter is $\bm{\phi}=\{\bm{\varphi},\bm{\omega}\}$.
(14) $\mathcal{L}_{\text{SAN}}(\bm{\phi};\bm{\psi},\bm{\eta},\bm{x})$ $=\mathcal{L}_{\text{GAN}}(\{\bm{\varphi},\bm{\omega}^{-}\};\bm{\psi},\bm{\eta},\bm{x})+\bm{\omega}^{\top}\left(h_{\bm{\varphi}^{-}}(\bm{x})-\mathbb{E}_{q_{\bm{\eta}}(\bm{z}|\bm{x})}[h_{\bm{\varphi}^{-}}(g_{\bm{\psi}}(\bm{z}))]\right),$
@@ Appendix B Editing procedure on guided generation
In Algorithm 1, we show the specific procedure for guided generation described in Section 3.4. Note that we apply a time-travel technique to the update rule in Section 3.4 as in (Lugmayr et al., 2022; Wang et al., 2022; Yu et al., 2023; He et al., 2024) to achieve better editing results. This technique adds noise after each gradient descent step to implicitly perform a multi-step optimization of minimizing the distance measuring function $\mathfrak{L}(\cdot,\bm{y})$ and lead to an improvement in the editing quality.
@@
0:  motion control signal $\bm{y}$, output of text encoder $\tau(\bm{c})$, estimator in latent diffusion $\epsilon_{\theta}(\cdot,t,\tau(\bm{c}))$, distance measuring function $\mathfrak{L}(\cdot;\bm{y})$, pre-defined parameter $\bar{\alpha}_{t}$, time-dependent step-size $\rho_{t}$, and the repeat times of time-travel of each step $\{r_{1},\cdots,r_{T}\}$.
@@
Appendix C Implementation Details
Training setup for stage 1 model: During the training, motion sequences are segmented into lengths of $L=64$. We use AdamW optimizer and batch size of 128. The hyperparameters in Eq. (4) and (6) are set as $\lambda_{\text{act}}=1.0$, $\lambda_{\text{reg}}=1.0\times 10^{-4}$, and $\lambda_{\text{adv}}=1.0\times 10^{-3}$. Our model $\{q_{\bm{\eta}},g_{\bm{\psi}},f_{\bm{\psi}}\}$ are trained with a simple multi-step learning late, the first 10,000 iterations with a learning rate of $2.0\times 10^{-4}$, and latter 5,000 with a learning rate of $2.0\times 10^{-5}$ in the case of HumanML3D dataset (Guo et al., 2022). Note that in Section E, we also conducted training and evaluation on a different dataset, the KIT-ML dataset (Plappert et al., 2016). For this dataset, the model was trained for the first 10,000 iterations with a learning rate of $5.0\times 10^{-5}$, followed by an additional 5,000 iterations with a reduced learning rate of $5.0\times 10^{-6}$. In addition, we replace the reconstruction loss in (4) with smooth $\ell_{1}$ loss (known as the special case of Huber loss) function and introduce a position enhancement term, following the technique in (Zhang et al., 2023b).
Training setup for stage 2 model: During the training, we use the AdamW optimizer with a batch size of 64. The model $\epsilon_{\theta}$ is trained for 10,000 epochs with a cosine annealing learning rate schedule starting at $1.0\times 10^{-4}$, including a warm-up phase. For inference, we adopt DDIM with 50 sampling steps and employ a trailing strategy. The classifier-free guidance scale $s$ in Eq. (19) is set to $s=11$ for HumanML3D and $s=7$ for the KIT-ML dataset. Notably, we observed in Section E that the performance is significantly influenced by the scale $s$.
Appendix D Evaluation metrics details